Angel finds $2.15 in Quarters and Nickels in her pocket. There were seven fewer quarters than nickels. How many nickels and quarters did she have?
Answer by ptaylor(2198)

Let x=number of nickels
then x-7=number of quarters
Now lets deal in pennies
5x+25(x-7)=215
5x+25x-175=215 add 175 to each side
30x-175+175=215+175
30x=390
x=13-------------number of nickels
x-7=13-7=6 -------------number of quarters
CK
5*13+25*6=215
65+150=215
215=215
Hope this helps--ptaylor